# Maximizing the efficiency of human feedback in AI alignment: a comparative analysis

### v1

@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a central paradigm for aligning machine learning systems with human preferences [christiano2017deep, ibarz2018reward, ziegler2019fine, ouyang2022training]. RLHF leverages human preference judgments to guide policy optimization, typically by fitting a reward model from pairwise comparisons of model outputs. The efficiency of this pipeline critically depends on how sample pairs are generated and how the resulting comparisons are aggregated. In the standard approach, pairs are generated at random and modeled using a Bradley–Terry framework, which assumes latent utility scores and fits them via maximum likelihood estimation. While statistically principled, this approach does not account for annotation cost, and its sample efficiency is limited under constrained human evaluation budgets. Notably, training the Bradley-Terry model with random pairs of samples hasn’t been contested since its first adoption by Christiano et al. [christiano2017deep]. This raises the following critical question:
@@ 3 Methodology
The Borda count method comes from the social choice theory and offers a simple yet effective positional voting rule where the score of each item is the number of wins it has at the end [mclean1995classics]. We apply this method in our tests with two different sampling strategies: either randomly ($N/2$ comparisons per round) or in a round-robin style manner known as Copeland’s method (Figure 2) in social choice theory [copeland1951reasonable, saari1996copeland] when the resources are unrestricted (requires N(N2)/2N(N-2)/2 comparisons per round). We refer the readers to Algorithm 2 for our implementation of the random sampling approach, with the Copeland variation being very similar.
@@
We also tried to implement the Quicksort approach based on [maystre2017quicksort] and assigning a Borda count score after each sorting round. The results for this technique were much poorer than the methods included in our experiments and, therefore, are omitted from the figures in our experiments later on, but it will be included, however, in our code-base release.
@@ 3.3 The Elo rating system
This method applies the Elo rating system Eqs. (2) & (3). We chose a simplified form instead of variations that are designed for specific applications (e.g, chess). The Elo can be updated either after each comparison or at the end of the round, with no differences in the results. The pairs can be either sampled randomly or using the Copeland method (Figure 2). We use Algorithm 16 to implement the random sampling approach, with the Copeland variation being very similar.
@@ 3.4 The Swiss tournament method
The Swiss tournament system involves pairing items based on their score [hua2017swiss]. In each round, every item is paired with the item with the closest running score. At the end of each round, all items update their Elo rating score and are paired again in the next round with other items [hua2017swiss]. We implement the Swiss tournament method using Algorithm 4.
@@
The Swiss tournament system – created for chess tournaments and widely applied and studied by game theory – involves pairing items based on their score [hua2017swiss]. In each round, every item is paired with the item with the closest running score. At the end of each round, all items update their Elo rating score and are paired again in the next round with other items [hua2017swiss]. While this method guarantees the best items come to the top and the worst go last, it leaves the middle point a little more clouded. Furthermore, when every item is originally unranked, applying it in its pure form can lead to discretised distributions of scores. We implement two variations of this method, one being the pure Swiss tournament method, with the other having random pairings in the beginning and then applying the Swiss method. Algorithm 5 describes the second approach, while setting the random rounds to zero, it regresses back to the pure Swiss tournament method. The reason for exploring random pairing rounds in the beginning is that the Elo system is path dependent, meaning that the order in which an item compares with others affects the final resulting rating. With all items having a more noisy initial rating, applying the Swiss method can further separate the items even faster. Items belonging in the middle of the distribution will have greater initial variance, thus leading to potentially better separation. Concerning the allocation of the annotation budget between random pairing rounds and Swiss pairing rounds, we examined all combinations in our simulations with the 50-50 rule, yielding, on average, better results. This finding is convenient for its applicability in RLHF, as the randomly sampled pairs can be pre-generated and therefore substantially reduce the computational cost required during the annotation session.
@@ 4 Experimental Evaluation
For our experiments, we evaluated the following techniques: the Bradley–Terry model estimator (Bradley–Terry), the Borda count method with random sampling (Borda–RNG), the Borda count method combined with Copeland’s sampling strategy (Borda–Copeland), random sampling followed by a Swiss tournament using the Elo rating system (Elo–RNG+Swiss), and a Swiss tournament with an information–gain–based pairing rule (Swiss–InfoGain). In addition, we explored several other approaches, but due to consistently poor performance, they are omitted from most of the figures reported later. These include Elo rating with purely random sampling (Elo–RNG), Elo rating combined with Copeland’s sampling strategy (Elo–Copeland), a pure Swiss tournament using Elo ratings (Swiss Tournament), and the Quicksort–based ranking methods of Maystre and Grossglauser [maystre2017quicksort], where Borda scores were assigned after sorting.
@@ 4.2 Comparing the method performance as $M$ increases
At this point, the question is intuitive: What happens for other budgets? Established practices expect the Bradley-Terry model to be the best for excessive annotation budgets but as we see some of the methods can be better for lower budgets. The question we try to answer here is what’s the tipping point considering again $N=100$ items. In this experiment we compare the following methods: Bradley-Terry, Borda-RND, Elo-RND and Elo RND+Swiss by conducting the simulations from budgets of 500 all the way to 20.000 comparisons. For statistical significance we use 100 random seeds and report the mean values in a bands with 95% confidence intervals. Note, the Copeland methods can’t be scaled to operate on the full range of 500 to 20.000 comparisons, so we only report the results in multiples of 4950 comparisons. As for the Swiss tournament method it operated on average 500 pairs across the 100 seeds and for the Swiss InfoGain method it needs on average less than 2500 pairs. The results are in Figure 6. There are two very interesting findings here. First, the Swiss InfoGain method outperformed everything else across most of the comparison budget range (considering $N=100$). This makes it very efficient for most annotation budgets. In fact it needs more 17,000 comparisons for Borda-Copeland method to clearly outperform it. For higher budgets the Borda-Copeland method remains the clear winner requiring a staggering 20,000 comparisons for the Bradley-Terry model to match it’s performance. These results indicate a very interesting design regime. For low to medium annotation budgets, the Swiss InfoGain algorithm is the most efficient and potent method to create item pairs, but after 16,000, the much simpler Borda count methods display the best performance.
@@ Appendix 0.A Algorithms
Algorithm 1 Bradley–Terry Estimator (with Gradient Descend)

### v2

@@ 1 Introduction
ReinforcementExtended version with appendix: https://arxiv.org/abs/2511.12796 Learning from Human Feedback (RLHF) has emerged as a central paradigm for aligning machine learning systems with human preferences [christiano2017deep, ibarz2018reward, ziegler2019fine, ouyang2022training]. RLHF leverages human preference judgments to guide policy optimization, typically by fitting a reward model from pairwise comparisons of model outputs. The efficiency of this pipeline critically depends on how sample pairs are generated and how the resulting comparisons are aggregated. In the standard approach, pairs are generated at random and modeled using a Bradley–Terry framework, which assumes latent utility scores and fits them via maximum likelihood estimation [christiano2017deep]. While statistically principled, this approach does not account for annotation cost, and its sample efficiency is limited under constrained human evaluation budgets [casper2023open, Wang2024ACS]. Notably, training the Bradley-Terry model with random pairs of samples hasn’t been contested since its first adoption by Christiano et al. [christiano2017deep]. This raises the following critical question:
@@ 3 Methodology
The Borda count method comes from the social choice theory and offers a simple yet effective positional voting rule where the score of each item is the number of wins it has at the end [mclean1995classics]. We apply this method in our tests with two different sampling strategies: either randomly ($N/2$ comparisons per round) or in a round-robin style manner known as Copeland’s method (Figure 2) in social choice theory [copeland1951reasonable, saari1996copeland] when the resources are unrestricted (requires N(N2)/2N(N-2)/2 comparisons per round). We refer the readers to Algorithm 2 for our implementation of the random sampling approach, with the Copeland variation being very similar. We also tried to implement the Quicksort approach based on [maystre2017quicksort] and assigning a Borda count score after each sorting round. The results for this technique were much poorer than the methods included in our experiments and, therefore, are omitted from the figures in our experiments later on, but it will be included, however, in our code-base release.
@@ 3.3 The Elo rating system
This method applies the Elo rating system Eqs. (2) & (3). We chose a simplified form instead of variations that are designed for specific applications (e.g, chess). The Elo can be updated either after each comparison or at the end of the round, with no differences in the results. The pairs can be either sampled randomly or using the Copeland method (Figure 2). Refer to Algorithm 3 for the random sampling approach, with the Copeland variation being very similar.
@@ 3.4 The Swiss tournament method
The Swiss tournament system involves pairing items based on their score [hua2017swiss] (Figure 3). In each round, every item is paired with the item with the closest running score. At the end of each round, all items update their Elo ratings and are paired again with other items in the next round [hua2017swiss] (Algorithm 4).
@@
The Swiss tournament system – created for chess tournaments and widely applied and studied by game theory – involves pairing items based on their score [hua2017swiss]. In each round, every item is paired with the item with the closest running score. At the end of each round, all items update their Elo ratings and are paired again with other items in the next round [hua2017swiss]. While this method guarantees the best items come to the top, and the worst go to the bottom, it leaves the middle point a little more clouded. Furthermore, when every item is originally unranked, applying it in its pure form can lead to discretised distributions of scores. We implement two variations of this method, one being the pure Swiss tournament method, with the other having random pairings in the beginning and then applying the Swiss method. Algorithm 5 describes the second approach, while setting the random rounds to zero, it regresses back to the pure Swiss tournament method. The reason for exploring random pairing rounds in the beginning is that the Elo system is path dependent, meaning that the order in which an item compares with others affects the final resulting rating. With all items having a more noisy initial rating, applying the Swiss method can further separate the items even faster. Items belonging in the middle of the distribution will have greater initial variance, thus leading to potentially better separation. Concerning the allocation of the annotation budget between random pairing rounds and Swiss pairing rounds, we examined all combinations in our simulations with the 50-50 rule, yielding, on average, better results. This finding is convenient for its applicability in RLHF, as the randomly sampled pairs can be pre-generated and therefore substantially reduce the computational cost required during the annotation session.
@@ 4 Experimental Evaluation
For our experiments, we evaluated the following techniques: i) the Bradley–Terry model estimator (Bradley–Terry), ii) the Borda count method with random sampling (Borda–RNG), iii) the Borda count method combined with Copeland’s sampling strategy (Borda–Copeland), iv) Elo rating with purely random sampling (Elo–RNG), v) Elo rating combined with Copeland’s sampling strategy (Elo–Copeland), vi) a pure Swiss tournament using Elo ratings (Swiss Tournament), vii) random sampling followed by a Swiss tournament using the Elo rating system (Elo–RNG+Swiss), and viii) our Swiss tournament system variation with an information–gain–based pairing rule (Swiss–InfoGain). In addition, we explored a few other approaches (e.g the Quicksort–based ranking methods of Maystre and Grossglauser [maystre2017quicksort]) that are omitted from our reports due to poor performance or similarity to the methods covered previously in this paper.
@@ 4.2 Comparing the method performance as $M$ increases
At this point, the question is intuitive: What happens for other budgets? Established practices expect the Bradley-Terry model to be the best for excessive annotation budgets, but as we see, some of the methods can be better for lower budgets. The question we try to answer here is what’s the tipping point, considering $N=100$ items again. In this experiment, we compare the following methods: Bradley-Terry, Borda-RND, Elo-RND, and Elo RND+Swiss by conducting the simulations from budgets of 500 all the way to 20.000 pairs. For statistical significance, we use 100 random seeds and report the mean values in bands with 95% confidence intervals. Note, the Copeland methods can’t be scaled to operate on the full range of 500 to 20.000 comparisons, so we only report the results in multiples of 4950 comparisons. As for the Swiss tournament method, it operated on average 500 pairs across the 100 seeds, and for the Swiss InfoGain method, it needs, on average, fewer than 2500 pairs. The results are in Figure 6. There are two very interesting findings here. First, the Swiss InfoGain method outperformed everything else across most of the comparison budget range (considering $N=100$). This makes it very efficient for most annotation budgets. In fact, it needs more than 15,000 pairs for the Borda-Copeland method to clearly outperform it (measuring to 8̃3% greater data efficiency). For higher budgets, the Borda-Copeland method remains the clear winner, requiring a staggering 20,000 pairs for the Bradley-Terry model to match its performance. Compared to Swiss InfoGain, the Bradley-Terry model requires more than 17,500 pairs to outperform it, measuring to 8̃6% greater data efficiency. These results point towards a very interesting design regime. For low to medium annotation budgets, the Swiss InfoGain algorithm is the most efficient and potent method to create item pairs, but for unrestricted annotation budgets, the Borda-Copeland performs best.
@@ 7 Conclusion
7.0.1 Acknowledgements
This work was supported in part by the EU Horizon projects with numbers 101092912 (MLSysOps) and 101160671 (DIGITISE).
@@ Appendix 0.A Algorithms
Algorithm 1 Bradley–Terry Estimator (with Gradient Descent)